OpenStudy (anonymous):

How can you sketch the graph y = |x - 2| ?

OpenStudy (anonymous):

Translate the graph of y = |x| two units left.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait, no... move it right, not left

OpenStudy (anonymous):

on the x axis, subtraction is to the right side and addition is to the left. its "backwards" from what you would expect

OpenStudy (anonymous):

cuz its absolute value

OpenStudy (anonymous):

no. thats just how translations work example y=x^2 and y=(x-2)^2 the second one is moved right 2 spaces

OpenStudy (anonymous):

Translations work that way because the equation is solved for y in terms of x. It's as if y is treating x as if it were 2 units smaller, so x needs to be bigger to have the same effect.

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

i never thought of it that way, thanks for the explanation @CliffSedge

OpenStudy (anonymous):

That's the only way it makes sense to me. :">